OpenStudy (anonymous):

Jordan wants to solve the following system using the elimination method: 2x + 3y = 10 x + y = 7 What number should the equation x + y = 7 be multiplied by to eliminate y? (4 points) –2 2 –3 3

OpenStudy (kropot72):

If you want to create an equation that has y eliminated, you need to make the coefficient of y in the equation x + y = 7 the same as the coefficient of y in the equation 2x + 3y = 10.

OpenStudy (kropot72):

What is the coefficient of y in the equation 2x + 3y = 10?

OpenStudy (anonymous):

3 right?

OpenStudy (kropot72):

Good work! 3 is correct. Therefore if you multiply each term in the equation x + y = 7 by 3, the coefficient of y will become the same as the coefficient of y in the equation 2x + 3y = 10.

OpenStudy (anonymous):

Thank you that helped a lot !

OpenStudy (kropot72):

You're welcome :)